let tan(x)=5/12 with x in quadrant 1,
find csc(2x)

Answer by lwsshak3(11628)

let tan(x)=5/12 with x in quadrant 1,
find csc(2x)
***
working with a (5-12-13) reference right triangle
sinx=5/13
cosx=12/13
csc(2x)=1/sin(2x)=1/(2sinxcosx)=1/(2*5/13*12/13)=1/(120/169)=169/120